OpenStudy (anonymous):

persamaan 3xkuadrat-4x = 5x mempunyai akar-akar x1 dan x2 dengan x1

OpenStudy (anonymous):

no english??

OpenStudy (anonymous):

\[3x ^{2}-4x=5x\]\[3x ^{2}-9x=0\]\[3x(x-9)=0\] \[x1\rightarrow 3x=0; \rightarrow x1=0\]\[x2\rightarrow x-9=0; x2=9\] understand what I mean ?

OpenStudy (anonymous):

\[3x ^{2}-4x=5x\]\[3x ^{2}-9x=0\]\[3x(x-3)=0\] \[x1\rightarrow 3x=0;\rightarrow x1=0\]\[x2\rightarrow x-3=0;\rightarrow x2=3\]

OpenStudy (anonymous):

akar-akar persamaan kuadrat 2x pangkat2 +3x-5=0 adalah..

OpenStudy (anonymous):

\[2x ^{2}+3x-5=0\] \[(2x+5)(x-1)=0\]\[2x+5=0; \rightarrow x1=-5/2\]\[x-1=0; \rightarrow x2=1\]